OpenStudy (anonymous):

State testing tomorrow. I'm studying and need help refreshing my memory so please, Explain. Sandra wrote the sequence below. 2,5,10,17,... Which equation represents the rule for finding the nth term of this sequence?

ganeshie8 (ganeshie8):

hint : perfect squares

ganeshie8 (ganeshie8):

1^2 + 1 = 2 2^2 + 1 = 5 3^2 + 1 = 10 4^2 + 1 = 17 . . n^2 + 1 is the nth term, eh ?

OpenStudy (anonymous):

ahhh... so it would be C. an = n^2 +1

ganeshie8 (ganeshie8):

Yes :)

OpenStudy (anonymous):

Could you help me with more? I really need to jog my memory! I forgot half of the math I learned and I feel like im going to fail this test.

ganeshie8 (ganeshie8):

sure :) solving them on ur own is the best thing for preparing exam... u can ask the ones you're stuck on...

OpenStudy (anonymous):

@ganeshie8 a team has twelve 15-year- old players and eight 16-year old players. The coach of the team is 43 years old. Which measure of the central tendency best represents the ages of the team, including the coach, why? for some reason most of the questions on the sample test is a little confusing because they don't include geometry.

ganeshie8 (ganeshie8):

this is very easy question to answer, if u knw one important thing about mean/median and the outliers

OpenStudy (anonymous):

what is outliers?

ganeshie8 (ganeshie8):

mean is NOTA a good measure if u have outliers in the data

ganeshie8 (ganeshie8):

yes the old coach is the outlier here

ganeshie8 (ganeshie8):

notice how majority of the ages are between 15 and 16 years only, but the coach age is 43 years... which lies very far outside from the major bulk of ages

ganeshie8 (ganeshie8):

Median is the best measure when u have outliers

ganeshie8 (ganeshie8):

since u have outlier in this data, Median is the best measure.

OpenStudy (anonymous):

Median, because it is not affected by the coaches age

ganeshie8 (ganeshie8):

thats the perfect way to put it !